if sin x = 3 sin( x + 2y)
then what is the value of tan( x+y ) + 2tan y

Grade:12

1 Answers

Hashan Hasalanka
6 years ago
Sin(x+y-y)=3sin(x+y+y)Sin(x+y)cos(y)-cos(x+y)sin(y) =3[sin(x+y)cos(y)+cos(x+y)sin(y)]Then devide it by cos(x+y)cos(y)You getTan(x+y)-tan(y) = 3[tan(x+y)+tan(y)]2tan(x+y)+4tan(y)=0SinceTan(x+y)+2tan(y)=0